How do i find the slope and y-intercept of 7x-y=4?
Answer by amoresroy(361)

How do i find the slope and y-intercept of 7x-y=4?
To find the slope and y-intercept of a line, express the equation in this form
y = mx + b
where m = slope
b = y-intercept
7x-y = 4
Subtract 7x to both sides
7x-7x-y = -7x + 4
-y = -7x +4
Multiply both sides by -1
y = 7x - 4
So the slope is 7 and the y-intercept is -4